At the equator, near the surface of the Earth, the magnetic field is approximately 50.0 μT northward, and the electric field is about 90.0 N/C downward in fair weather. Find the gravitational, electric, and magnetic forces on an electron in this environment, assuming that the electron has an instantaneous velocity of 9.00 ✕ 106 m/s directed to the east.

gravitational force     N directed  ---Select--- upward downward
electric force     N directed  ---Select--- upward downward eastward westward
magnetic force     N directed  ---Select--- upward downward eastward westward
Answer #1

a)

Fg = mg = 9.11* 10^-31* 9.8 = 8.928* 10^-30 N

downwards

b)

Fe = qE = 1.6* 10^-19* 90 = 1.44* 10^-17 N

upwrads

c)

Fm = qvB = 1.6* 10^-19* 9* 10^6* 50* 10^-6 = 7.2* 10^-17 N

downward